# Stochastic Gradient Langevin Unlearning

### v1

@@ 3.2 Unlearning guarantees with PABI analysis
Feldman et al. (2018) provide the following metric-aware bound of the Rényi divergence between two $c$-CNI processes, which is simplified by Altschuler & Talwar (2022b).
Lemma 3.4 (Metric-aware PABI bound (Feldman et al., 2018), simplified by Altschuler & Talwar (2022b) in Proposition 2.10).
@@
In light of Lemma 3.4, if we can characterize the initial $W_{\infty}$ distance, the Rényi divergence between two $c$-CNI processes can be bounded. Note that the PNSGD update is indeed a $c$-CNI with $c<1$ under strong convexity assumption. We further provide the following two lemmas to track the $W_{\infty}$ distance along the learning process.
@@
Sketch of proof. First note that the PNSGD update leads to a $(1-\eta m)$-CNI process when $\eta\leq 1/L$ for any mini-batch sequence. Recall that $y_{k}^{j}$ is the unlearning process at epoch $k$ at iteration $j$, starting from $y_{0}^{0}=x_{T}^{0}=\mathcal{M}(\mathcal{D})$. Consider the “adjacent” process $y_{k}^{j,\prime}$ starting from $y_{0}^{0,\prime}=x_{T}^{0,\prime}=\mathcal{M}(\mathcal{D}^{\prime})$ but still fine-tune on $\mathcal{D}^{\prime}$ so that $y_{k}^{j},y_{k}^{j,\prime}$ only differ in their initialization. Now, consider three distributions: $\nu_{\mathcal{D}|\mathcal{B}},\nu_{T|\mathcal{B}}^{0},\rho_{K|\mathcal{B}}^{0}$ are the stationary distribution for the learning processes, learning process at epoch $T$ and unlearning process at epoch $K$ respectively. Similarly, consider the “adjacent” processes that learn on $\mathcal{D}^{\prime}$ and still unlearn on $\mathcal{D}^{\prime}$ (see Figure 1 for the illustration). Denote distributions $\nu_{\mathcal{D}^{\prime}|\mathcal{B}},\nu_{T|\mathcal{B}}^{0,\prime},\rho_{K|\mathcal{B}}^{0,\prime}$ for these processes similarly. Note that our goal is to bound $d_{\alpha}(\rho_{K|\mathcal{B}}^{0},\nu_{T|\mathcal{B}}^{0,\prime})$ for the RU guarantee. By weak triangle inequality (Mironov, 2017), we can upper bound it in terms of $d_{2\alpha}(\rho_{K|\mathcal{B}}^{0},\nu_{\mathcal{D}^{\prime}|\mathcal{B}})$ and $d_{2\alpha}(\nu_{\mathcal{D}^{\prime}|\mathcal{B}},\nu_{T|\mathcal{B}}^{0,\prime})$, which are the $\varepsilon_{2}$ and $\varepsilon_{1}$ terms in Theorem 3.2 respectively. For $d_{\alpha}(\nu_{\mathcal{D}^{\prime}|\mathcal{B}},\nu_{T|\mathcal{B}}^{0,\prime})$, we leverage the naive $2R$ bound for the $W_{\infty}$ distance between $\nu_{\mathcal{D}^{\prime}|\mathcal{B}},\nu_{0|\mathcal{B}}^{0,\prime}$ and applying Lemma 3.4 leads to the desired result. For $d_{\alpha}(\rho_{K|\mathcal{B}}^{0},\nu_{\mathcal{D}^{\prime}|\mathcal{B}})$, by triangle inequality of $W_{\infty}$ and Lemma 3.53.6 one can show that the $W_{\infty}$ between $\rho_{0|\mathcal{B}}^{0},\nu_{\mathcal{D}^{\prime}|\mathcal{B}}$ is bounded by $Z_{\mathcal{B}}$ in Theorem 3.2. Further applying Lemma 3.4 again completes the proof.
@@
Our proof only relies on the bounded gradient difference $\|\nabla f(x;\mathbf{d})-\nabla f(x;\mathbf{d}^{\prime})\|\leq 2M$ $\forall x\in\mathbb{R}^{d}$ and $\forall\mathbf{d},\mathbf{d}^{\prime}\in\mathcal{X}$ hence $M$-Lipchitz assumption can be replaced. In practice, we can leverage the gradient clipping along with a $\ell_{2}$ regularization for the convex objective function.
The convergent case. In practice, one often requires the training epoch $T$ to be sufficiently large so that the model is “well-trained”, where a similar assumption is made in the prior unlearning literature (Guo et al., 2020; Sekhari et al., 2021). Under this assumption, we can further simplify Theorem 3.2 into the following corollary.
@@
$\varepsilon\leq\frac{\alpha Z_{\mathcal{B}}^{2}}{2\eta\sigma^{2}}c^{2Kn/b},Z_{\mathcal{B}}=\min\left(\frac{1}{1-c^{n/b}}\frac{2\eta M}{b},2R\right).$
For simplicity, the rest of the discussion on our stochastic gradient Langevin unlearning will based on the training convergent assumption.
From Corollary 3.8 one can observe that a smaller $b$ leads to a better decaying rate ($c^{2Kn/b}$) but also a potentially worse initial distance $Z_{\mathcal{B}}=O(1/((1-c^{n/b})b))$. In general, choosing a smaller $b$ still leads to less epoch for achieving the desired privacy loss. In practice, choosing $b$ too small (e.g., $b=1$) can not only degrade the utility but also incur instability (i.e., large variance) of the convergent distribution $\nu_{\mathcal{D}|\mathcal{B}}$, as $\nu_{\mathcal{D}|\mathcal{B}}$ depends on the design of mini-batches $\mathcal{B}$. One should choose a moderate $b$ to balance between privacy and utility, which is the unique privacy-utility-complexity trade-off with respect to $b$ revealed by our analysis.
Computational benefit compared to retraining. In the view of Corollary 3.8 or Lemma 3.4, it is not hard to see that a smaller initial $W_{\infty}$ distance leads to fewer PNSGD (un)learning epochs for being $\varepsilon$-close to a target distribution $\nu_{\mathcal{D}^{\prime}|\mathcal{B}}$ in terms of Rényi difference $d_{\alpha}$. For stochastic gradient Langevin unlearning, we have provided a uniform upper bound $Z_{\mathcal{B}}=O(\eta M/((1-c^{n/b})b))$ of such initial $W_{\infty}$ distance in Lemma 3.5. On the other hand, even if both $\nu_{0},\nu_{\mathcal{D}}$ are both Gaussian with identical various and mean difference norm of $\Omega(1)$, we have $W_{\infty}(\nu_{0},\nu_{\mathcal{D}^{\prime}|\mathcal{B}})=\Omega(1)$ for retraining from scratch. Our results show that a larger mini-batch size $b$ leads to more significant complexity savings compared to retraining. As we discussed above, one should choose a moderate size $b$ to balance between privacy and utility. In our experiment, we show that for commonly used mini-batch sizes (i.e., $b\geq 32$), our stochastic gradient Langevin unlearning is still much more efficient in complexity compared to retraining.
Improved bound with randomized $\mathcal{B}$. So far our results are based on a fixed (worst-case) mini-batch sequence $\mathcal{B}$. One can improve the privacy bound in Corollary 3.8 by taking the randomness of $\mathcal{B}$ into account under a non-adaptive unlearning setting. That is, the unlearning request is independent of the mini-batch sequence $\mathcal{B}$. See also our discussion in the related work. We leverage Lemma 4.1 in Ye & Shokri (2022), which is an application of joint convexity of KL divergence. By combining with Corollary 3.8, we arrive at the following improved bound.
@@
where $Z_{\mathcal{B}}$ is the bound described in Lemma 3.5.
It slightly improves the privacy bound of Corollary 3.8 by the average case of $\mathcal{B}$ instead of the worst case.
@@
There are several remarks for Corollary 3.10. First, the privacy loss now only decays linearly instead of exponentially as opposed to the strongly convex case. Second, $Z_{\mathcal{B}}$ now can grow linearly in training epoch $T$. As a result, the computational benefit of our approach compared to retraining may vanish in this scenario. We conjecture a better analysis is needed beyond strong convexity.

### v2

@@ 2.1 Converting Initial $W_{\infty}$ Distance to Final Rényi Divergence Bound
Lemma 2.6 (Metric-aware privacy amplification by iteration bound [20], simplified by [16] in Proposition 2.10).
@@
Roughly speaking, Lemma 2.6 shows that if we have the $W_{\infty}$ distance of initial distributions of two PNSGD processes on the same dataset, we have the corresponding Rényi difference bound after $K$ iterations. The projection set diameter $2R$ is a default upper bound for $W_{\infty}$ distance if we do not care about the initial distributions as the case studied in  [16] for mixing time analysis. In the unlearning scenario, the initial distributions of the unlearning processes, denoted by $\nu_{\mathcal{D}|\mathcal{B}},\nu_{\mathcal{D}^{\prime}|\mathcal{B}}$, are much more relevant. We can show a much tighter bound by analyzing $W_{\infty}(\nu_{\mathcal{D}|\mathcal{B}},\nu_{\mathcal{D}^{\prime}|\mathcal{B}})$ along the adjacent PNSGD learning processes.
3 Certified Unlearning Guaranatee for PNSGD
We start with introducing the (un)learning process with PNSGD with a cyclic mini-batch strategy (Algorithm 1). Note that this mini-batch strategy is not only commonly used for practical DP-SGD implementations in privacy libraries [21], but also in theoretical analysis for DP guarantees [22]. For the learning mechanism $\mathcal{M}$, we optimize the objective function with PNSGD on dataset $\mathcal{D}$ (line 3-8 in Algorithm 1). $\eta,\sigma^{2}>0$ are hyperparameters of step size and noise variance respectively. The initialization $\nu_{0}$ is an arbitrary distribution supported on $\mathcal{C}_{R}$ if not specified. For the unlearning mechanism $\mathcal{U}$, we fine-tune the current parameter with PNSGD on the updated dataset $\mathcal{D}^{\prime}$ subject to the unlearning request (line 10-15 in Algorithm 1) with $y_{0}^{0}=x_{T}^{0}=\mathcal{M}(\mathcal{D})$. For the rest of the paper, we denote $\nu_{t}^{j},\rho_{k}^{j}$ as the probability density of $x_{t}^{j},y_{k}^{j}$ respectively. Furthermore, we denote =$\mathcal{B}=\{\mathcal{B}^{j}\}_{j=0}^{n/b-1}$ the minibatch sequence described in Algorithm 1, where $b$ is the step size and we assume $n$ is divided by $b$ throughout the paper for simplicity [^2] When $n$ is not divided by $b$, we can simply drop the last $n-\lfloor n/b\rfloor b$ points.. We use $\nu_{\cdot|\mathcal{B}}$ to denote the conditional distribution of $\nu_{\cdot}$ given $\mathcal{B}$.
3.1 Certified Unlearning Guarantees
Now we introduce the certified unlearning guarantees for PNSGD and the corresponding analysis illustrated in Figure 1. We first prove that for any fixed mini-batch sequence $\mathcal{B}$, the limiting distribution $\nu_{\mathcal{D}|\mathcal{B}}$ of the learning process exists, is unique, and stationery. The proof is deferred to Appendix E and is based on applying the results in [23] to establish the ergodicity of the learning process $x_{t}^{0}$.
Theorem 3.1.
Suppose that the closed convex set $\mathcal{C}\subset\mathbb{R}^{d}$ is bounded with $\mathcal{C}$ having a positive Lebesgue measure and that $\nabla f(\cdot;\mathbf{d}_{i}):\mathcal{C}\to\mathbb{R}^{d}$ is continuous for all $i\in[n]$. The Markov chain $\{x_{t}:=x_{t}^{0}\}$ in Algorithm 1 for any fixed mini-batch sequence $\mathcal{B}$ admits a unique invariant probability measure $\nu_{\mathcal{D}|\mathcal{B}}$ on the Borel $\sigma$-algebra of $\mathcal{C}$. Furthermore, for any $x\in\mathcal{C}$, the distribution of $x_{t}$ conditioned on $x_{0}=x$ converges weakly to $\nu_{\mathcal{D}|\mathcal{B}}$ as $t\to\infty$, where $\nu_{\mathcal{D}|\mathcal{B}}$ is the conditional distribution of $\nu_{\mathcal{D}}$ given $\mathcal{B}$.
Suppose the training epoch $T$ is large enough so that the model is well-trained for now, which means $\mathcal{M}(\mathcal{D})\sim\nu_{\mathcal{D}|\mathcal{B}}=\rho_{0|\mathcal{B}}^{0}$ and our target “retraining distribution” is $\nu_{\mathcal{D}^{\prime}|\mathcal{B}}$. Our goal is then to upper bound the Rényi difference $d_{\alpha}(\rho_{K|\mathcal{B}}^{0},\nu_{\mathcal{D}^{\prime}|\mathcal{B}})$ after $K$ unlearning epochs. In the case of insufficient training, the privacy loss is $d_{\alpha}(\rho_{K|\mathcal{B}}^{0},\nu_{T|\mathcal{B}}^{0,\prime})$ where $\mathcal{M}(\mathcal{D}^{\prime})\sim\nu_{T|\mathcal{B}}^{0,\prime}$ for $T$ training epochs. It can be upper bounded in terms of $d_{\alpha}(\rho_{K|\mathcal{B}}^{0},\nu_{\mathcal{D}^{\prime}|\mathcal{B}})$ and $d_{\alpha}(\nu_{T|\mathcal{B}}^{0,\prime},\nu_{\mathcal{D}^{\prime}|\mathcal{B}})$ via weak triangle inequality of Rényi divergence, which is provided in Theorem 3.2 below. We later provide a better bound by considering the randomness of $\mathcal{B}$.
Theorem 3.2 (RU guarantee of PNSGD unlearning, fixed $\mathcal{B}$).
Assume $\forall\mathbf{d}\in\mathcal{X}$, $f(x;\mathbf{d})$ is $L$-smooth, $M$-Lipchitz and $m$-strongly convex in $x$. Let the learning and unlearning processes follow Algorithm 1 with $y_{0}^{0}=x_{T}^{0}=\mathcal{M}(\mathcal{D})$. Given any fixed mini-batch sequence $\mathcal{B}$, for any $\alpha>1$, let $\eta\leq\frac{1}{L}$, the output of the $K^{th}$ unlearning iteration satisfies $(\alpha,\varepsilon)$-RU for any adjacent dataset $\mathcal{D},\mathcal{D}^{\prime}$, where
$\varepsilon\leq\frac{\alpha-1/2}{\alpha-1}\left(\varepsilon_{1}(2\alpha)+\varepsilon_{2}(2\alpha)\right),\varepsilon_{1}(\alpha)=\frac{\alpha(2R)^{2}}{2\eta\sigma^{2}}c^{2Tn/b},\;\varepsilon_{2}(\alpha)=\frac{\alpha Z_{\mathcal{B}}^{2}}{2\eta\sigma^{2}}c^{2Kn/b},$
$Z_{\mathcal{B}}=W_{\infty}(\rho^{0}_{K|\mathcal{B}},\nu_{\mathcal{D}^{\prime}|\mathcal{B}})\leq 2Rc^{Tn/b}+\min\left(\frac{1-c^{Tn/b}}{1-c^{n/b}}\frac{2\eta M}{b},2R\right),$
and $c=1-\eta m$.
As we explained earlier, we need non-trivial $W_{\infty}$ bounds for adjacent PNSGD processes to obtain better unlearning guarantees when applying Lemma 2.6. We provide such results below and the proofs are deferred to Appendix L and M respectively.
@@ 3.1 Certified Unlearning Guarantees
Remark 3.5.
In [15], the authors used the default projection set diameter $2R$ as the $W_{\infty}$ distance upper bound for their DP results, see equations (3.5) and (5.2) therein. However, it yields a vacuous bound as an unlearning guarantee compared to retraining from scratch. Note that our tighter bound for $W_{\infty}$ distance is also useful for deriving later sequential unlearning guarantees compared to the prior work based on the analysis of Langevein dynamics [11]. Interestingly, this improved result can also be utilized for tightening the DP guarantee in [15] and make it more practically useful, as $2R$ can be very large in practice, which may be of independent interest.
@@
Sketch of proof. First note that the PNSGD update leads to a $(1-\eta m)$-CNI process when $\eta\leq 1/L$ for any mini-batch sequence. Recall that $y_{k}^{j}$ is the unlearning process at epoch $k$ at iteration $j$, starting from $y_{0}^{0}=x_{T}^{0}=\mathcal{M}(\mathcal{D})$. Consider the “adjacent” process $y_{k}^{j,\prime}$ starting from $y_{0}^{0,\prime}=x_{T}^{0,\prime}=\mathcal{M}(\mathcal{D}^{\prime})$ but still fine-tune on $\mathcal{D}^{\prime}$ so that $y_{k}^{j},y_{k}^{j,\prime}$ only differ in their initialization. Now, consider three distributions: $\nu_{\mathcal{D}|\mathcal{B}},\nu_{T|\mathcal{B}}^{0},\rho_{K|\mathcal{B}}^{0}$ are the stationary distribution for the learning processes, learning process at epoch $T$ and unlearning process at epoch $K$ respectively. Similarly, consider the “adjacent” processes that learn on $\mathcal{D}^{\prime}$ and still unlearn on $\mathcal{D}^{\prime}$ (see Figure 1 for the illustration). Denote distributions $\nu_{\mathcal{D}^{\prime}|\mathcal{B}},\nu_{T|\mathcal{B}}^{0,\prime},\rho_{K|\mathcal{B}}^{0,\prime}$ for these processes similarly. Note that our goal is to bound $d_{\alpha}(\rho_{K|\mathcal{B}}^{0},\nu_{T|\mathcal{B}}^{0,\prime})$ for the RU guarantee. By weak triangle inequality [19], we can upper bound it in terms of $d_{2\alpha}(\rho_{K|\mathcal{B}}^{0},\nu_{\mathcal{D}^{\prime}|\mathcal{B}})$ and $d_{2\alpha}(\nu_{\mathcal{D}^{\prime}|\mathcal{B}},\nu_{T|\mathcal{B}}^{0,\prime})$, which are the $\varepsilon_{2}$ and $\varepsilon_{1}$ terms in Theorem 3.2 respectively. For $d_{\alpha}(\nu_{\mathcal{D}^{\prime}|\mathcal{B}},\nu_{T|\mathcal{B}}^{0,\prime})$, we leverage the naive $2R$ bound for the $W_{\infty}$ distance between $\nu_{\mathcal{D}^{\prime}|\mathcal{B}},\nu_{0|\mathcal{B}}^{0,\prime}$ and applying Lemma 2.6 leads to the desired result. For $d_{\alpha}(\rho_{K|\mathcal{B}}^{0},\nu_{\mathcal{D}^{\prime}|\mathcal{B}})$, by triangle inequality of $W_{\infty}$ and Lemma 3.33.4 one can show that the $W_{\infty}$ between $\rho_{0|\mathcal{B}}^{0},\nu_{\mathcal{D}^{\prime}|\mathcal{B}}$ is bounded by $Z_{\mathcal{B}}$ in Theorem 3.2. Further applying Lemma 2.6 again completes the proof.
@@
Our proof only relies on the bounded gradient difference $\|\nabla f(x;\mathbf{d})-\nabla f(x;\mathbf{d}^{\prime})\|\leq 2M$ $\forall x\in\mathbb{R}^{d}$ and $\forall\mathbf{d},\mathbf{d}^{\prime}\in\mathcal{X}$ hence $M$-Lipchitz assumption can be replaced. In practice, we can leverage the gradient clipping along with a $\ell_{2}$ regularization for the convex objective function [22].
The convergent case. In practice, one often requires the model to be “well-trained”, where a similar assumption is made in the prior unlearning literature [7, 8]. Under this assumption, we can further simplify Theorem 3.2 into the following corollary.
@@
$\varepsilon\leq\frac{\alpha Z_{\mathcal{B}}^{2}}{2\eta\sigma^{2}}c^{2Kn/b},\quad Z_{\mathcal{B}}=\min\left(\frac{1}{1-c^{n/b}}\frac{2\eta M}{b},2R\right).$
For simplicity, the rest of the discussion on our PNSGD unlearning will based on the well-trained assumption. From Corollary 3.7 one can observe that a smaller $b$ leads to a better decaying rate ($c^{2Kn/b}$) but also a potentially worse initial distance $Z_{\mathcal{B}}=O(1/((1-c^{n/b})b))$. In general, choosing a smaller $b$ still leads to less epoch for achieving the desired privacy loss. In practice, choosing $b$ too small (e.g., $b=1$) can not only degrade the utility but also incur instability (i.e., large variance) of the convergent distribution $\nu_{\mathcal{D}|\mathcal{B}}$, as $\nu_{\mathcal{D}|\mathcal{B}}$ depends on the design of mini-batches $\mathcal{B}$. One should choose a moderate $b$ to balance between privacy and utility, which is the unique privacy-utility-complexity trade-off with respect to $b$ revealed by our analysis.
Computational benefit compared to retraining. In the view of Corollary 3.7 or Lemma 2.6, it is not hard to see that a smaller initial $W_{\infty}$ distance leads to fewer PNSGD (un)learning epochs for being $\varepsilon$-close to a target distribution $\nu_{\mathcal{D}^{\prime}|\mathcal{B}}$ in terms of Rényi difference $d_{\alpha}$. For PNSGD unlearning, we have provided a uniform upper bound $Z_{\mathcal{B}}=O(\eta M/((1-c^{n/b})b))$ of such initial $W_{\infty}$ distance in Lemma 3.3. On the other hand, even if both $\nu_{0},\nu_{\mathcal{D}}$ are both Gaussian with identical various and mean difference norm of $\Omega(1)$, we have $W_{\infty}(\nu_{0},\nu_{\mathcal{D}^{\prime}|\mathcal{B}})=\Omega(1)$ for retraining from scratch. Our results show that a larger mini-batch size $b$ leads to more significant complexity savings compared to retraining. As we discussed above, one should choose a moderate size $b$ to balance between privacy and utility. In our experiment, we show that for commonly used mini-batch sizes (i.e., $b\geq 32$), our PNSGD unlearning is still much more efficient in complexity compared to retraining.
Improved bound with randomized $\mathcal{B}$. So far our results are based on a fixed (worst-case) mini-batch sequence $\mathcal{B}$. One can improve the privacy bound in Corollary 3.7 by taking the randomness of $\mathcal{B}$ into account under a non-adaptive unlearning setting. That is, the unlearning request is independent of the mini-batch sequence $\mathcal{B}$. See also our discussion in the related work. By taking the average of the bound in Corollary 3.7 in conjunction with an application of joint convexity of KL divergence [22], we can derive an improved guarantee beyond the worst-case of $\mathcal{B}$.
@@
Under the same setting as of Theorem 3.2 but with random mini-batch sequences described in Algorithm 1. Then for any $\alpha>1$, and $\eta\leq\frac{1}{L}$, the output of the $K^{th}$ unlearning iteration satisfies $(\alpha,\varepsilon)$-RU with $c=1-\eta m$, where $\varepsilon\leq\frac{1}{\alpha-1}\log\left(\mathbb{E}_{\mathcal{B}}\exp\left(\frac{\alpha(\alpha-1)Z_{\mathcal{B}}^{2}}{2\eta\sigma^{2}}c^{2Kn/b}\right)\right)$ and $Z_{\mathcal{B}}$ is the bound described in Lemma 3.3.
@@
Discussion on utility bound. One can leverage the utility analysis in section 5 of [24] to derive the utility guarantee for the full batch setting $b=n$. We relegate the proof to Appendix K.
Proposition 3.9.
Under the same setting as Corollary 3.7 with $b=n$, $\eta\leq\frac{m}{2L^{2}}$ and assume the minimizer of any $f_{\mathcal{D}}$ is in the relative interior of $\mathcal{C}_{R}\subseteq\mathbb{R}^{d}$, for any given adjacent dataset pair $\mathcal{D},\mathcal{D}^{\prime}$ the output of the $K^{th}$ unlearning iteration $y_{K}^{0}$ satisfies
$\mathbb{E}[f_{\mathcal{D}^{\prime}}(y_{K}^{0})-\inf_{x\in\mathcal{C}_{R}}f_{\mathcal{D}^{\prime}}(x)]\leq Mc^{K}\min(\frac{1}{1-c}\frac{2\eta M}{n},2R)+\frac{2Ld\sigma^{2}}{m}.$ (2)
Note that a similar analysis applies to both the mini-batch (i.e., $b\leq n$) and non-convergent case (i.e., Theorem 3.2) but the result is more complicated. We leave the rigorous analysis as future work. An important remark is that the second term $\frac{2Ld\sigma^{2}}{m}$ is the excess risk of $\nu_{\mathcal{D}^{\prime}}$, which is controlled by the noise scale $\sigma$. This presents the privacy-utility trade-off as demonstrated in the DP scenario in [24].
@@
There are several remarks for Corollary 3.10. First, the privacy loss now only decays linearly instead of exponentially as opposed to the strongly convex case. Second, $Z_{\mathcal{B}}$ now can grow linearly in training epoch $T$. As a result, the computational benefit of our approach compared to retraining may vanish for large $T$ such that $\frac{2\eta MT}{b}\geq 2R$. Nevertheless, the computational benefit against retraining persists for moderate $T$ such that $\frac{2\eta MT}{b}<2R$. This condition can be met if the model learns reasonably well with moderate $T$ and the projection diameter $2R$ is not set to be extremely small. For example, with $12R=10,b=128,\eta=$ 1 and $M=1$, any training epoch $T<640$ will lead to $\frac{2\eta MT}{b}<2R$. Still, we conjecture a better analysis is needed beyond strong convexity.